### v5

@@ Abstract
SPSysML validation is survey- and application-based. First, a robotic system for an assisted living project was developed. The integrity of simulated and physical parts of the system is improved using SPSysML-based quantitative evaluation. Thus, more system components are shared between the simulated and physical setups. The system was deployed on the physical robot and two simulators based on the Robot Operating System (ROS) or ROS2. SPSysML was used by a third-party developer and was assessed by him and other practitioners in a survey.
@@ 1 Introduction
Simulation is widely used in state-of-the-art cyber-physical systems (CPS) development procedures. Recent papers refer to Model-in-the-loop (MIL) [1], Software-in-the-loop (SIL) [2], Hardware-in-the-loop (HIL) [3] and Rapid Control Prototyping (RCP) [4] techniques. They are used selectively or are composed in sequence, e.g. verification and validation steps in a systems development approach called V-model [5]. Each of the techniques requires a simulation of the system parts. Some systems utilise the Digital Twin (DT) [6, 7] concept. They employ an accurate simulation of a system part for, e.g. Swapping malfunctioned parts, Energy consumption analysis, Technology integration, or Real-time monitoring.
@@
Developing reliable systems, especially complex ones like robot systems, requires comprehensive unit and integration testing. Some CPS parts can be tested with simulated hardware only; therefore, additional parts that do not comprise the operational system are required (e.g. human simulator in a social robot case). Comprehensive testing is complex in test case specification and time-consuming in test implementation and execution. From this perspective, software reusability is a key to fast development of complex, reliable systems.
@@
To reach these goals, we propose a Domain-Specific Language (DSL) named Simulated-Physical System Modeling Language (SPSysML). SPSysML models the requirements and the SPSys, in particular, SPSys consisting of simulated parts (used either in the system development or in the operational setup as a DT/simulator). Based on SPSysML, we propose a requirement-based SPSys structuring method for enhancing requirement readability in simulated and physical system parts. To measure the inter-embodiment integrity of the system, we propose quantitative design evaluation factors for SPSys designs. The factors are used during system development to maximise the shared controller, minimise the variety of the system parts and improve simulated-physical integrity. Our goal originates from the conclusion of [11]: The digital models are mainly used to examine product performance(…). However, how to optimise the use of those models to enhance the design process and design collaboration still needs to be investigated.

### v6

@@ Abstract
SPSysML validation is survey- and application-based. First, we develop a robotic system for an assisted living project. As a result of the SPSysML application, we observed an integrity improvement between the simulated and physical parts of the system. Thus, more system components are shared between the simulated and physical setups. The system was deployed on the physical robot and two simulators based on ROS and ROS2. Additionally, we share a questionnaire for SPSysML assessment. The feedback that we already received is published in this article.
@@
{graphicalabstract}[Uncaptioned image]
{highlights}
Goal: We propose an evaluation method of integrity between the simulated and physical embodiment of the system. The assessment is done in various scopes, e.g. per pair of Digital Twins (DT) and its physical counterpart— Physical Twin (PT), system-wide, or one of many system setups.
Method: We propose a Domain-Specific Language based on Systems Modeling Language (SysML). The Simulated-Physical Systems Modeling Language (SPSysML) defines the taxonomy of considered CPS consisting of at least a physical or simulated part. Based on SPSysML concepts, we define quantitative factors and a requirement-based system structuring method, which enhances requirement analysis and allows DT to perceive exogenous actions in the simulated world.
Result: SPSysML is used to develop a robotic system for the INCARE project. In subsequent iterations of the system’s design process, the simulated-physical integrity of the system is improved, and a more significant fraction of the system components is shared between its simulated and physical embodiments. The designed system was deployed on the physical robot and two simulators. System setups are based on Robot Operating System (ROS) or ROS2. Therefore, we argue that SPSysML is neither specific for a control system framework nor a robot simulator. SPSysML was used by a third-party developer and was assessed by him and other practitioners in a survey.
Conclusion: SPSysML enables the design of SPSys (including DT and CPS) and evaluating the design in their multi-setup system development. The system features improved integrity between simulated and physical parts in its setups. The proposed systematic requirement-based system structuring also enhances the traceability of system requirements allocation to simulated and physical system parts.
@@ 1 Introduction
Simulation is widely used in state-of-the-art cyber-physical systems (CPS) development procedures. Recent papers refer to Model-in-the-loop (MIL) [46], Software-in-the-loop (SIL) [12], Hardware-in-the-loop (HIL) [39] and Rapid Control Prototyping (RCP) [1] techniques. They are used selectively or are composed in sequence, e.g. verification and validation steps in a systems development approach called V-model [37]. Each of the techniques requires a simulation of the system parts. Some systems utilise the Digital Twin (DT) [33, 47] concept. They employ an accurate simulation of a system part for, e.g. Swapping malfunctioned parts, Energy consumption analysis, Technology integration, or Real-time monitoring. Making robot control systems trustworthy is one of the biggest challenges [9]. Application of DT-based testing to robot system development is a promising and popular solution to this problem. However, DT integration to robotic systems raises, e.g. synchronisation issues [23] and software integrity issues tackled in this article.
@@
Developing reliable systems, especially complex ones like robot systems, requires well-designed system architectures [3, 10], often defined as Domain-specification language (DSL) and comprehensive unit and integration testing. Some CPS parts can be tested with simulated hardware only; therefore, additional parts that do not comprise the operational system are required (e.g. human simulator in a social robot case). Comprehensive testing is complex in test case specification and time-consuming in test implementation and execution. From this perspective and from issues in robot application in Industry 4.0 [25], software reusability is a key to the fast development of complex, reliable systems.
@@
To reach these goals, we propose a DSL named Simulated-Physical System Modeling Language (SPSysML). It is built upon Systems Modeling Language (SysML), the standard and popular language used in the systems engineering domain [62]. SPSysML models the requirements and the SPSys, in particular, SPSys consisting of simulated parts (used either in the system development or in the operational setup as a DT/simulator). Based on SPSysML, we propose a requirement-based SPSys structuring method for enhancing requirement traceability in simulated and physical system parts. To measure the inter-embodiment integrity of the system, we propose quantitative design evaluation factors for SPSys designs. The factors are used during system development to maximise the shared controller, minimise the variety of the system parts and improve simulated-physical integrity. Our goal originates from the conclusion of [32]: The digital models are mainly used to examine product performance(…). However, how to optimise the use of those models to enhance the design process and design collaboration still needs to be investigated.